# Inference Scaling fLaws: The Limits of LLM Resampling with Imperfect Verifiers

### v2

@@ Abstract
Recent research has generated hope that inference scaling could allow weaker language models to match or exceed the accuracy of stronger models, such as by repeatedly sampling solutions to a coding problem until it passes unit tests. The central thesis of this paper is that there is no free lunch for inference scaling: indefinite accuracy improvement through resampling can only be realized if the “verifier” (in this case, a set of unit tests) is perfect. When the verifier is imperfect, as it almost always is in domains such as reasoning or coding (for example, unit tests have imperfect coverage), there is a nonzero probability of false positives: incorrect solutions that pass the verifier. Resampling cannot decrease this probability, so it imposes an upper bound to the accuracy of resampling-based inference scaling even with an infinite compute budget. We find that there is a very strong correlation between the model’s single-sample accuracy (i.e. accuracy without unit tests) and its false positive rate on coding benchmarks HumanEval and MBPP, whose unit tests have limited coverage. Therefore, no amount of inference scaling of weaker models can enable them to match the single-sample accuracy of a sufficiently strong model (Fig. 1(a)). When we consider that false positives have a negative utility compared to abstaining from producing a solution, it bends the inference scaling curve further downward. Empirically, we find that the optimal number of samples can be less than 10 under realistic assumptions (Fig. 1(b)). Finally, we show that beyond accuracy, false positives may have other undesirable qualities, such as poor adherence to coding style conventions.
(a) Single-sample accuracy vs. resampling limits on HumanEval+. The x-axis shows single-sample accuracy on HumanEval+ (which contains comprehensive unit tests), while the y-axis shows the highest achievable accuracy when resampling with an infinite compute budget, using HumanEval’s more limited unit tests as verifiers. Weaker models (models with lower single-sample accuracy) produce false positive solutions at higher rates. Models below the cutoff line are unable to match GPT-4o through resampling, as GPT-4o’s Pass@1 exceeds the accuracy of such a model even when conditioned on its solutions passing the unit tests. Results on MBPP+ follow a similar pattern (Fig. 3).
(b) The cost of false positives limits the reward of resampling. False positives w.r.t. an imperfect verifier (HumanEval unit tests) incur a "cost" (e.g., subtle bugs in code), while correct answers provide a benefit. The reward (y-axis) depends on this cost-benefit ratio (C/B-Ratio). Problem instances that require more attempts tend to be harder, hence more susceptible to false positives. Thus, even with zero computational cost, for realistic cost-benefit ratios, the optimal number of samples $K$ is finite and very low.
@@ 1 Introduction
Scaling the amount of compute used during inference is a promising way to improve LLM performance. Techniques include reasoning [69, 71, 53], reflecting on model outputs to revise candidate solutions [55, 80], and compositions of these and other atomic techniques [51, 70].
Inference scaling through resampling stands out for its simplicity and broad applicability. It works by generating many candidate outputs until one is satisfactory, based on feedback from a verifier [58, 49, 7, 22, 36]. Unlike techniques such as majority voting where gains from inference scaling quickly plateau (Table 1), resampling has given rise to the hope of usefully scaling inference compute by many orders of magnitude.
@@
Review of inference scaling techniques and their limitations (Section 2). We review papers on inference scaling, categorizing the primary techniques and listing their domain-specific applications and known limitations.
Demonstration of generalization gap (Section 3). We provide empirical evidence on two benchmarks, HumanEval+ and MBPP+ [42], showing that the apparent gains from resampling with imperfect verifiers are unlikely to translate into real-world performance. Despite achieving comparable results to stronger models on standard unit tests, less capable models suffer from a larger generalization gap—producing incorrect solutions that fail the extended test suite (false positives) at higher rates than stronger models.
@@
Empirical analysis to understand the limitations of inference scaling with imperfect verifiers (Section 4). We examine how introducing a cost (negative utility) for returning false positives impacts the optimal number of resampling attempts on HumanEval+. We find that even with an infinite inference budget, the optimal number of samples is often finite and very low (e.g., $K\leq 5$ in Fig. 4). Hence, resampling quickly reaches a point of diminishing returns without bridging the performance gap for smaller models. If the cost of an incorrect solution is higher than the benefit of a correct solution, the optimal $K$ can be zero — the risk of a false positive for a weak model is high enough that it is effectively useless (Fig. 4). In Section A.3 we present a theoretical model that complements the findings in this section.
Evidence that this affects code quality beyond correctness (Section 5). We show that the reliance on imperfect verifiers not only affects the functional correctness but also overall quality of the generated code. We evaluate candidate solutions on HumanEval+ based on various readability metrics such as adherence to naming conventions (that we specify in the prompt) like snake_case and camelCase, line-level commenting, and guidelines regarding the maximum line length and number of lines in function implementations. We find that false positive solutions are lower quality across all models and metrics when compared to true positive solutions. While other aspects of code quality such as simplicity and modularity are harder to test automatically, we speculate that the same pattern holds for those properties as well.
@@
Our findings have two additional implications. First, they show the importance of building highly accurate verifiers. This goal might benefit from treating verification technology as a specialized subfield with its own metrics and benchmarks. This is especially true for training-time uses: models trained with feedback from imperfect verifiers may learn to exploit weaknesses in the verifier rather than solve tasks robustly, potentially leading to safety concerns [35, 2].
Second, the use of imperfect verifiers as the ground truth for evaluation is flawed. We used HumanEval+ and MBPP+ for evaluation; the gaps we identify would have been invisible if we had used HumanEval and MBPP both as verifiers and as benchmarks. While the limitations of these benchmarks for measuring absolute performance are well known [78, 42], our results show that they might result in misleading comparisons between models as well.
@@ 2 Scaling inference compute with verifiers
Technique Description Limitations Reasoning Applying structured logical steps to improve output quality [69] Performance gains are unreliable and domain-dependent [44, 60, 34] Can increase the likelihood to produce harmful or undesirable output [54] Critique Self-evaluating and refining initial outputs [55] Improvements from self-refinement worsens performance for tasks with high model uncertainty [59, 27, 63] Self-refinement amplifies self-bias [73] Fusion Combining multiple samples into one output [51] Unreliable performance improvements that vary by task [51] Ranking Scores and ranks the best samples from multiple candidates [14, 22, 41, 40, 45, 26, 33, 53, 57, 64, 13] Doesn’t scale with sample budget [7] Underperform compared to other methods [77] Majority Voting Using consensus among multiple samples to determine final output[68, 37, 67] Hurts performance on hard tasks and non-monotonous scaling under task heterogeneity [10] Sample inefficient for queries with many answer possibilities [67] Limited applicability for tasks with non-discrete answers Oracle Verification Leverages ground-truth evaluator for free until correct solution is found [72, 19] Not available for most domains Imperfect Verification
@@
Scores and accepts or rejects candidate solutions [77, 16, 75, 21, 30] Bigger generalization gap for weaker models (Section 3) Optimal number of samples is finite and low (Section 4) Low code quality of false positives (Section 5)
Table 1 provides an overview of the main techniques for scaling inference compute with LLMs. Some methods such as majority voting [68, 10] or resampling using verifiers [7, 72] generate many candidate solutions and then select one. Other methods such as reasoning [69] and critique [55, 47] refine a single solution. In practice, these methods can be combined in flexible ways and the distinction between them is not always clear (Section A.1.1). Note that our notion of inference scaling excludes methods such as those used to train OpenAI’s o1 series of models, since we are only looking at improvements during inference time to available language models, rather than training improvements.
All these methods except verifier-based resampling are known to have important limitations that cast doubt on how much scaling is truly possible, as summarized in Table 1. Resampling using verifiers has a different control flow than other methods, which gives it an intuitive appeal (Fig. 2): we can potentially regenerate solutions indefinitely until one is correct. This enthusiasm around resampling is partly driven by the empirically observed inference scaling laws, which suggest that the fraction of tasks for which we find at least one correct solution scales predictably with the number of samples over multiple orders of magnitude [7].
However, the usefulness of this depends on the availability of a capable verifier [16]. In some settings, we may have an oracle verifier, such as a proof checker, that does not suffer from false positives — that is, if the proof checker verifies the proof, it is guaranteed to be correct. False negatives of the verifier (including nontermination under a fixed compute budget) are less of a problem, as one can simply generate more samples until a true positive is found. It is possible that every correct solution is a false negative of the verifier, but it is unclear if this is a problem that arises in practice.
But in other settings such as coding and reasoning, we only have imperfect verifiers such as unit tests or LM judges, which suffer from false positives: incorrect solutions that nonetheless pass the verifier. In these settings, we don’t have easy methods to guarantee the correctness of generated solutions at inference time. As a result, we cannot distinguish between false positives and true positives simply by increasing the compute budget. We survey papers that use verifiers in Table 2.
Paper Verifier Category Verifier Type Domain Verifier Implementation Chen et al. [9] Imperfect Unit tests Coding Checks agreement of tests and samples Shinn et al. [55] Imperfect LM-as-judge,
@@
Yao et al. [75] Imperfect LM-as-judge Planning LLM evaluates reasoning steps First et al. [19] Oracle Proof checker Math Proof checker Thakur et al. [61] Oracle Proof checker Math Proof checker Yang et al. [74] Oracle Proof checker Math Proof checker Wang et al. [66] Oracle Proof checker Math Proof checker Azerbayev et al. [4] Oracle Proof checker Math Proof checker, Majority voting Huang et al. [28] Oracle Proof checker Math Proof checker Xin et al. [72] Oracle Proof checker Math Proof checker Brown et al. [7] Oracle Proof checker Math Proof checker Davis et al. [16] Imperfect LM-as-judge QA, Math LLM judges correctness of generations Hassid et al. [22] Imperfect Unit tests Coding Unit tests Zhang et al. [77] Imperfect Generative RM Math Verification as part of the model output Zhuge et al. [81] Imperfect Agent-as-judge Agents, Coding Agents evaluate outputs of other agents Kapoor et al. [32] Imperfect Unit tests Coding Unit tests Saad-Falcon et al. [51] Imperfect LM-as-judge Coding, Reasoning LLM judges correctness of generations Liang et al. [39] Imperfect Program-of-thought Coding, Math Checks CoT against generated PoT Cook et al. [15] Imperfect LM-as-judge Instruction-following LLM checks answer against generated checklists Gundawar et al. [21] Imperfect Agent-as-judge Travel planning Pre-defined constraints verified by critic agents
@@
In this paper, we investigate the effect of scaling inference compute with access to imperfect verifiers.
@@ 3 Repeated sampling with weaker models leads to worse generalizability
In computer programming tasks, unit tests are commonly employed as verifiers to assess the correctness of candidate solutions generated by language models. While unit tests are practical and efficient, they often suffer from imperfect test coverage, leading to false positives where incorrect solutions pass the tests [20]. This affects many benchmarks such as HumanEval [11], APPS [24], or MBPP [3]. This imperfection raises the question: Do less capable models produce false positives—implementations that pass the standard unit tests but fail the comprehensive ones—at a higher rate than stronger models?
@@
Experimental setup. To investigate this, we conducted experiments on two widely used coding benchmarks: HumanEval+ and MBPP+. MBPP consists of simple programming tasks designed to evaluate the basic coding abilities of models [3]. HumanEval+ and MBPP+ are extensions of the original HumanEval and MBPP benchmarks [42] and contain additional hidden test cases to assess correctness beyond the unit tests included in the original benchmarks.
@@
We evaluated multiple models of varying capabilities, including weaker and stronger models, generating at least 50 samples for each model and benchmark task. We used the standard unit tests provided with the benchmarks as imperfect verifiers to filter the candidate solutions. To assess the generalization gap, we then evaluated solutions that passed the original benchmark test sets on the more comprehensive hidden test cases (see Section A.5 for details). These tests are extensive, and we assume that solutions that pass the full set of tests are correct. (If this assumption is not true, the generalization gaps that we reveal only grow bigger.)
@@
Findings. Weaker models exhibit a higher probability of producing false positives compared to stronger models (Fig. 3). The false positive probability scales inversely with the true capability. This linear relationship holds with remarkable consistency across models of various model families, including Cohere’s Command models, GPT-4o, and the Llama 3.1 family. This suggests that while weaker models appear to perform well on standard benchmarks through increased sampling, they fail to generalize effectively and, importantly, they generalize worse than more capable models. They tend to generate fragile solutions that exploit the limitations of the unit tests. We speculate that this is because weaker models’ “true understanding” of the programming tasks is worse.
@@
The empirical results reinforce a core insight. Suppose $\mathbb{P}_{strong}(\text{Correct})>\mathbb{P}_{weak}(\text{Correct}|\text{Pass Verifier})$. That is, the single-sample accuracy of a strong model exceeds that of a weaker model, even conditioned on the weaker model passing the base unit tests. Then the weaker model cannot match the performance of a single invocation of the stronger model, no matter how big the compute budget for the weaker model. In Fig. 3, this is shown by a horizontal line. No model below the line can match the performance of GPT-4o through resampling.
@@
The effect is largely driven by a subset of the tasks where the unit tests are poor. When we limit the analysis to these tasks, the relationship is even more pronounced Section A.2.
@@
(a) HumanEval+
(b) MBPP+
Figure 3: Generalization gap with infinite compute budget. We show the relationship between the accuracy of individual samples (x-axis) and the achievable accuracy given an infinite compute budget and limited unit tests (y-axis; note that it starts at 0.5). We evaluate performance on the extended test suites of HumanEval+ and MBPP+, using the unit tests from the original benchmarks as verifiers. For both benchmarks, the trend is that less capable models are more likely to generate false positives than stronger models. In Section A.2, we show our results with the full y-axis as well as upper (lower) bounds on the conditional accuracy accounting for tasks for which we did not observe any solutions passing the unit tests.

### v3

# The Limits of Inference Scaling Through Resampling

@@ Abstract
Recent research has generated hope that inference scaling, such as resampling solutions until they pass verifiers like unit tests, could allow weaker models to match stronger ones. Beyond inference, this approach also enables training reasoning models, where data is curated using rejection sampling against a verifier. However, we show that this approach is fundamentally limited when verifiers are imperfect and have a non-zero probability of producing false positives. Resampling cannot decrease this probability, so it imposes an upper bound to the accuracy of resampling-based inference scaling, regardless of compute budget. Our analysis shows that there is a strong correlation between the model’s single-sample accuracy and its false positive rate on HumanEval and MBPP, whose unit tests have limited coverage. Therefore, no amount of inference scaling of weaker models can enable them to match the single-sample accuracy of a sufficiently strong model. Empirical results show that optimal sampling attempts are often fewer than 10, as the negative utility of false positives outweighs benefits, bending inference scaling curves downward. Finally, false positives may have other undesirable qualities, like poor adherence to coding style conventions.
Refer to caption Single-sample accuracy vs. resampling limits on HumanEval+. The x-axis shows single-sample accuracy on HumanEval+ (which contains comprehensive unit tests), while the y-axis shows the highest achievable accuracy when resampling with an infinite compute budget, using HumanEval’s more limited unit tests as verifiers. Weaker models (models with lower single-sample accuracy) produce false positive solutions at higher rates. Models below the cutoff line are unable to match GPT-4o through resampling, as GPT-4o’s Pass@1 exceeds the accuracy of such a model even when conditioned on its solutions passing the unit tests. Results on MBPP+ follow a similar pattern (Figure˜3).
Refer to caption The cost of false positives limits the reward of resampling. False positives w.r.t. an imperfect verifier (HumanEval unit tests) incur a "cost" (e.g., subtle bugs in code), while correct answers provide a benefit. The reward (y-axis) depends on this cost-benefit ratio (C/B-Ratio). Problem instances that require more attempts tend to be harder, hence more susceptible to false positives. Thus, even with zero computational cost, for realistic cost-benefit ratios, the optimal number of samples $K$ is finite and very low.
@@ 1 Introduction
Scaling the amount of compute used during inference is a promising way to improve LLM performance. Techniques include reasoning (Wei et al., 2023; Wu et al., 2024; Setlur et al., 2024), reflecting on model outputs to revise candidate solutions (Shinn et al., 2023; Zhong et al., 2024), and compositions of these and other atomic techniques (Saad-Falcon et al., 2024; Welleck et al., 2024).
Inference scaling through resampling stands out for its simplicity and broad applicability. It works by generating many candidate outputs until one is satisfactory, based on feedback from a verifier (Song et al., 2024; Qin et al., 2024; Brown et al., 2024; Hassid et al., 2024; Li et al., 2024a). Unlike techniques such as majority voting where gains from inference scaling quickly plateau (Table˜1), resampling has given rise to the hope of usefully scaling inference compute by many orders of magnitude.
@@
Review of inference scaling techniques and their limitations (Section˜2). We review papers on inference scaling, categorizing the primary techniques and listing their domain-specific applications and known limitations.
Demonstration of generalization gap (Section˜3). We provide empirical evidence on two benchmarks, HumanEval+ and MBPP+ (Liu et al., 2023b), showing that the apparent gains from resampling with imperfect verifiers are unlikely to translate into real-world performance. Despite achieving comparable results to stronger models on standard unit tests, less capable models suffer from a larger generalization gap—producing incorrect solutions that fail the extended test suite (false positives) at higher rates than stronger models.
@@
Empirical analysis to understand the limitations of inference scaling with imperfect verifiers (Section˜4). We examine how introducing a cost (negative utility) for returning false positives impacts the optimal number of resampling attempts on HumanEval+. We find that even with an infinite inference budget, the optimal number of samples is often finite and very low (e.g., $K\leq 5$ in Figure˜4). Hence, resampling quickly reaches a point of diminishing returns without bridging the performance gap for smaller models. If the cost of an incorrect solution is higher than the benefit of a correct solution, the optimal $K$ can be zero — the risk of a false positive for a weak model is high enough that it is effectively useless (Figure˜4). In Appendix˜C we present a theoretical model that complements the findings in this section.
Evidence that this affects code quality beyond correctness (Section˜5). We show that the reliance on imperfect verifiers not only affects the functional correctness but also overall quality of the generated code. We evaluate candidate solutions on HumanEval+ based on various readability metrics such as adherence to naming conventions (that we specify in the prompt) like snake_case and camelCase, line-level commenting, and guidelines regarding the maximum line length and number of lines in function implementations. We find that false positive solutions are lower quality across all models and metrics when compared to true positive solutions. While other aspects of code quality such as simplicity and modularity are harder to test automatically, we speculate that the same pattern holds for those properties as well.
@@
Our findings have three additional implications. First, they show the importance of building highly accurate verifiers. This goal might benefit from treating verification technology as a specialized subfield with its own metrics and benchmarks. This is especially true for training-time uses: models trained with feedback from imperfect verifiers may learn to exploit weaknesses in the verifier rather than solve tasks robustly, potentially leading to safety concerns (Krakovna et al., 2020; Amodei et al., 2016).
Second, the use of imperfect verifiers as the ground truth for evaluation is flawed. We used HumanEval+ and MBPP+ for evaluation; the gaps we identify would have been invisible if we had used HumanEval and MBPP both as verifiers and as benchmarks. While the limitations of these benchmarks for measuring absolute performance are well known (Zhang et al., 2024b; Liu et al., 2023b), our results show that they might result in misleading comparisons between models as well.
Third, our findings highlight limitations in resampling-based data curation for reasoning models. Reasoning models rely on datasets curated through rejection sampling against verifiers. When these verifiers are imperfect, the curated datasets risk including mislabeled examples, which incurs a cost on model performance. This introduces a bottleneck: without stronger base models or highly accurate verifiers, the gains from resampling-based data curation to train reasoning models are likely limited.
@@ 2 Scaling inference compute with verifiers
Technique Description Limitations Ranking Scores and ranks the best samples from multiple candidates (Cobbe et al., 2021; Hassid et al., 2024; Liu et al., 2023a; Lightman et al., 2023; Liu et al., 2023c; Hosseini et al., 2024b; Kirchner et al., 2024; Setlur et al., 2024; Snell et al., 2024; Vacareanu et al., 2024; Chen et al., 2024c) Doesn’t scale with sample budget (Brown et al., 2024) Underperform compared to other methods (Zhang et al., 2024a) Majority Voting Using consensus among multiple samples to determine final output(Wang et al., 2023b; Li et al., 2024b; Wang et al., 2024b) Hurts performance on hard tasks and non-monotonous scaling under task heterogeneity (Chen et al., 2024a) Sample inefficient for queries with many answer possibilities (Wang et al., 2024b) Limited applicability for tasks with non-discrete answers Oracle
Verification
Leverages ground-truth evaluator for free until correct solution is found (Xin et al., 2024; First et al., 2023) Not available for most domains
[…]
Verification
@@
Scores and accepts or rejects candidate solutions (Zhang et al., 2024a; Davis et al., 2024; Yao et al., 2023; Gundawar et al., 2024; Kambhampati et al., 2024) Bigger generalization gap for weaker models (Section˜3) Optimal number of samples is finite and low (Section˜4) Low code quality of false positives (Section˜5)
Table˜1 provides an overview of the main techniques for scaling inference compute with LLMs. Some methods such as majority voting (Wang et al., 2023b; Chen et al., 2024a) or resampling using verifiers (Brown et al., 2024; Xin et al., 2024) generate many candidate solutions and then select one. Other methods such as reasoning (Wei et al., 2023) and critique (Shinn et al., 2023; Madaan et al., 2023) refine a single solution. In practice, these methods can be combined in flexible ways and the distinction between them is not always clear (Section˜A.1). Note that our notion of inference scaling excludes methods such as those used to train OpenAI’s o1 series of models, since we are only looking at improvements during inference time to available language models, rather than training improvements.
All these methods except verifier-based resampling are known to have important limitations that cast doubt on how much scaling is truly possible, as summarized in Table˜1. Resampling using verifiers has a different control flow than other methods, which gives it an intuitive appeal (Figure˜2): we can potentially regenerate solutions indefinitely until one is correct. This enthusiasm around resampling is partly driven by the empirically observed inference scaling laws, which suggest that the fraction of tasks for which we find at least one correct solution scales predictably with the number of samples over multiple orders of magnitude (Brown et al., 2024).
However, the usefulness of this depends on the availability of a capable verifier (Davis et al., 2024). In some settings, we may have an oracle verifier, such as a proof checker, that does not suffer from false positives — that is, if the proof checker verifies the proof, it is guaranteed to be correct. False negatives of the verifier (including nontermination under a fixed compute budget) are less of a problem, as one can simply generate more samples until a true positive is found. It is possible that every correct solution is a false negative of the verifier, but it is unclear if this is a problem that arises in practice.
[…]
@@
[…]
@@
But in other settings such as coding and reasoning, we only have imperfect verifiers such as unit tests or LM judges, which suffer from false positives: incorrect solutions that nonetheless pass the verifier. In these settings, we don’t have easy methods to guarantee the correctness of generated solutions at inference time. As a result, we cannot distinguish between false positives and true positives simply by increasing the compute budget. We survey papers that use verifiers in Table˜2.
In this paper, we investigate the effect of scaling inference compute with access to imperfect verifiers. We distinguish ranking from verification: ranking selects the best among multiple candidates, while verification accepts or rejects each candidate independently against a correctness criterion.
@@ 3 Repeated sampling with weaker models leads to worse generalizability
In computer programming tasks, unit tests are commonly employed as verifiers to assess the correctness of candidate solutions generated by language models. While unit tests are practical and efficient, they often suffer from imperfect test coverage, leading to false positives where incorrect solutions pass the tests (Gulwani et al., 2017). This affects many benchmarks such as HumanEval (Chen et al., 2021), APPS (Hendrycks et al., 2021), or MBPP (Austin et al., 2021). This imperfection raises the question: Do less capable models produce false positives—implementations that pass the standard unit tests but fail the comprehensive ones—at a higher rate than stronger models?
@@
Experimental setup. To investigate this, we conducted experiments on two widely used coding benchmarks: HumanEval+ and MBPP+. MBPP consists of simple programming tasks designed to evaluate the basic coding abilities of models (Austin et al., 2021). HumanEval+ and MBPP+ are extensions of the original HumanEval and MBPP benchmarks (Liu et al., 2023b) and contain additional hidden test cases to assess correctness beyond the unit tests included in the original benchmarks.
@@
We evaluated models of varying capabilities, including weak and stronger models, generating at least 200 samples per task and model on HumanEval+ (1,000 for Command Light, to minimize tasks without any passing solutions) and 50 samples per task on MBPP+. To assess the generalization gap, we then evaluated solutions that passed the original benchmark test sets on the more comprehensive hidden test cases (see Appendix˜D for details). These tests are extensive, and we assume that solutions that pass the full set of tests are correct. (If this assumption is not true, the generalization gaps that we reveal only grow bigger.)
@@
Findings. Weaker models exhibit a higher probability of producing false positives compared to stronger models (Figure˜3). This probability scales inversely with the true capability. This linear relationship holds with remarkable consistency across models of various families, including Cohere’s Command models, GPT-4o, and the Llama-3.1 family. This suggests that while weaker models appear to perform well on standard benchmarks through increased sampling, they fail to generalize effectively and, importantly, they generalize worse than more capable models. They tend to generate fragile solutions that exploit the limitations of the unit tests. We speculate that this is because weaker models’ “true understanding” of the programming tasks is worse.
@@
The empirical results reinforce a core insight. Suppose $\mathbb{P}_{strong}(\text{Correct})>\mathbb{P}_{weak}(\text{Correct}|\text{Pass Verifier})$. That is, the single-sample accuracy of a strong model exceeds that of a weaker model, even conditioned on the weaker model passing the base unit tests. Then the weaker model cannot match the performance of a single invocation of the stronger model, no matter how big the compute budget for the weaker model. In Figure˜3, this is shown by a horizontal line. No model below the line can match the performance of GPT-4o through resampling.
HumanEval+
MBPP+
Figure 3: Generalization gap with infinite compute budget. We show the relationship between the accuracy of individual samples (x-axis) and the achievable accuracy given an infinite compute budget and limited unit tests (y-axis; note that it starts at 0.5). We evaluate performance on the extended test suites of HumanEval+ and MBPP+, using the unit tests from the original benchmarks as verifiers. For both benchmarks, the trend is that less capable models are more likely to generate false positives than stronger models. In Appendix˜B, we show our results with the full y-axis as well as upper (lower) bounds on the conditional accuracy accounting for tasks for which we did not observe any solutions passing the unit tests.
@@
The effect is largely driven by a subset of the tasks where the unit tests are poor. When we limit the analysis to these tasks, the relationship is even more pronounced Appendix˜B.